# CARE: Confidence-Aware Regression Estimation of building density fine-tuning EO Foundation Models

### v1

@@ IV Evaluation and Results of CARE
In Table I, we evaluate CARE and examine both the average error and the median error. The median error of CARE is low, $0.00184$; this is desirable as it shows that the difference between the absolute error of the regression output and the estimated metric $(1-\text{confidence})$ is small. We also evaluate the regression output result using the MSE, like in (3), in Table I. We run experiments using $n$-sample tests, similar to [1], where we use $n$ training samples per region for fine-tuning. Stratified sampling is performed, as in the PhilEO Bench, and the regions are: Denmark, East Africa, Egypt, Guinea, Europe, Ghana, Israel, Japan, Nigeria, North America, Senegal, South America, Tanzania and Uganda [1, 5]. The performance of CARE, for $n=5000$, is examined in Table I.
For the evaluation of CARE, we also compare our model with other models. In Table I, we compare the results we obtain when we use Negative Log Likelihood (NLL) [17] and the model Gaussian-Output NLL. The percentage improvement of CARE compared to Gaussian-Output NLL is $30.34\%$ for the MSE, and $23.87\%$ for the average error. Moreover, we also compare CARE with the model Error-Sorting Confidence [6] in Table I. The percentage improvement of our model compared to Error-Sorting Confidence is $2.40\%$ for the MSE, and $92.83\%$ for the mean error. The percentage improvement of CARE compared to the model Absolute-Error Confidence is $0.91\%$ for the MSE, and $10.39\%$ for the average error. The Absolute-Error Confidence is also an ablation study, using only the first term in (4). These results demonstrate the efficacy of our algorithm and the superiority of CARE in outputting a reliable confidence metric for pixel-wise regression tasks, leading to both trustworthy and accurate model predictions.
@@
We thus have compared CARE to the Gaussian-Output NLL model in Table I, where the latter minimizes the Gaussian negative logarithmic likelihood loss. Hence, the MSE loss for the regression error achieves higher accuracy (i.e. $0.00326$ compared to $0.00468$), and this is one of our main targets.
For the evaluation of CARE, in addition to the average and median absolute errors, we also compute the correlation between the confidence and the error in the regression. The correlation is $0.62090$. For the model Gaussian-Output NLL, it is $0.56696$. Here, the percentage improvement is $9.51\%$.
We also compare CARE to the ensembles method with $M=\{3,1\}$ members [18]. By combining the Gaussian distributions from the ensemble members using a heteroscedastic Gaussian log-likelihood loss, this method estimates the epistemic uncertainty. For $n=1000$, for $M=3$, the MSE and mean error are $0.00685$ and $0.01306$, respectively, while for $M=1$, these are $0.00696$ and $0.01353$. For CARE, in Table II, these are $0.0039$ and $0.0138$. Here, because of (2) and (3), the percentage improvement in MSE of our model CARE, compared to the $M=3$ ensembles method [18], is $43.1\%$.
Furthermore, for the evaluation of CARE, we also examine in Table I the MSE $20\%$ or $10\%$ which is a threshold, i.e. $\zeta$, for detecting instances where the model simply does not know the correct result (i.e. pixel-wise regression value) from the available input data, e.g. due to lack of spectral information or resolution. In such cases, models might choose to abstain from providing an answer and should be able to output “None of the above” for the segmentation regression result value. A building density of $20\%$ with an estimated error of ±$\pm 10\%$ leads to the predicted density of $20\%$ being not useful, i.e. not accurate. On the contrary, a building density of $80\%$ with an estimated error of ±$\pm 1\%$ leads to the predicted density of $80\%$ being useful and accurate. A prediction by the model with a high confidence indicates high reliability and trust for this particular prediction. Also, the features for high building density might not be clear in multi-spectral optical EO data. This is due to epistemic uncertainty, as it is induced by the lack of detail in the measurement. This is why a plausible output set that is an Open Set, is needed. Epistemic uncertainty is systematic, caused by lack of knowledge, and can be reduced by learning the characteristics of the quantity (e.g., high building density) using additional information (for example, in-situ measurements). Moreover, aleatoric uncertainty is statistical and related to randomness, and the sample not being a typical example of the quantity (e.g., high or low building density).
@@
We evaluate the proposed model CARE in Table II using $n$-sample tests, for $n$ from $50$ to $10000$ samples. We examine both the MSE for the accuracy of the regression, i.e. (3) in (2), and the average error for the correctness of the confidence metric, i.e. of $c_{i}$ in (4). Here, we observe that for $n=7500$ samples, the MSE is $0.00316$, while for $n=500$, the MSE is $0.0043$, i.e. the performance of the model improves when we have more labeled data. In addition, for the accuracy of the confidence metric of our model, the median error in Table II is $0.00190$ for $n=7500$ samples, while the median error is $0.0088$ for $n=500$. The performance of the model CARE, for confidence quantification, improves when we have more labeled data, i.e. $n=7500$ samples compared to $n=500$.
@@
For the evaluation of our model CARE, we also examine in Table II the MSE $20\%$ or $10\%$, similar to in Table I. We observe that the performance of the proposed model CARE for estimating the building density, i.e. how close the buildings are in the Sentinel-2 multi-spectral images, improves when the model has the ability to abstain in cases of low confidence. These results demonstrate the effectiveness of our algorithm in providing a reliable confidence metric for regression that leads to trustworthy and more accurate model predictions.
@@
In Table III, we evaluate the model Error-Sorting Confidence [6] on the Sentinel-2 PhilEO Bench building density estimation downstream task [1], to compare these results with the results of the proposed model CARE in Table II. Error-Sorting Confidence is used in the fine-tuning stage, i.e. when starting from and using a trained EO FM [1, 3]. We observe that for $n=100$ training samples per region, the percentage improvement of CARE compared to Error-Sorting Confidence is $8.93\%$ for the MSE, and $73.47\%$ for the median error.
For the evaluation of CARE, we also examine qualitative results, in Figs. 1-3. For the error in the prediction, i.e. in (e), the variation in the regression output is, for example, because we underestimate in (b) building density. In (f), the predicted uncertainty is $(1-\text{confidence})$, i.e. using (d). In the colorbar in (f), the maximum value is smaller than $1$, and this is desirable, i.e. we have less highlighted and emphasized yellow color.
@@ V Conclusion
In this paper, we have proposed the model CARE for pixel-wise regression tasks. Our model CARE computes and assigns a confidence metric to the output results of deep neural networks that solve regression problems. The focal point of this work has been confidence quantification and assessment of regression algorithms that are based on EO Foundation Models applied on downstream tasks, i.e. building density estimation from satellite Sentinel-2 data. Using the proposed model CARE, we are able to estimate confidence as an unbiased predictor of the error rate. In this work, we have evaluated the proposed model CARE using the average absolute error between the absolute error of the regression output and the metric $(1-\text{confidence})$. We have demonstrated that the model CARE is effective and is able to both predict building density and estimate confidence. In addition, we have also compared our model with other baseline models and we have shown that the proposed model CARE outperforms other approaches.

### v2

@@ IV-A Evaluation of CARE using absolute error in Table I and comparison to other models
For the evaluation of CARE, we also compare our model with other models. In Table I, we compare the results we obtain when we use Negative Log Likelihood (NLL) [17] and the model Gaussian-Output NLL. The percentage improvement of CARE compared to Gaussian-Output NLL is $30.34\%$ for the MSE, and $23.87\%$ for the average error. We observe that according to these results, for the use case of building density estimation, regarding the model Gaussian-Output NLL, an interesting open research question is how useful the normal distribution assumption is. Moreover, we also compare CARE with the model Error-Sorting Confidence [6] in Table I. The percentage improvement of our model compared to Error-Sorting Confidence is $2.40\%$ for the MSE, and $92.83\%$ for the mean absolute error.
The percentage improvement of CARE compared to the model Absolute-Error Confidence is $0.91\%$ for the MSE, and $10.39\%$ for the average error. The Absolute-Error Confidence is also an ablation study, using only the first term in (4). These results demonstrate the efficacy of our algorithm and the superiority of CARE in outputting a reliable confidence metric for pixel-wise regression tasks, leading to both trustworthy and accurate model predictions.
@@
We therefore have compared CARE to the Gaussian-Output NLL model in Table I, where the latter minimizes the Gaussian negative logarithmic likelihood loss. Hence, the MSE loss for the regression error achieves higher accuracy (i.e. $0.00326$ compared to $0.00468$), and this is one of our main targets. For the evaluation of CARE, in addition to the average and median absolute errors, we also compute the correlation between the confidence and the error in the regression. The correlation is $0.62090$. For the model Gaussian-Output NLL, it is $0.56696$. Here, the percentage improvement is $9.51\%$.
We also compare the proposed model CARE to the ensembles method with $M=\{3,1\}$ members [18]. By combining the Gaussian distributions from the ensemble members using a heteroscedastic Gaussian log-likelihood loss, this method estimates the epistemic uncertainty. For $n=1000$, for $M=3$, the MSE and mean absolute error are $0.00685$ and $0.01306$, respectively, while for $M=1$, these are $0.00696$ and $0.01353$. For CARE, in Table II, these are $0.0039$ and $0.0138$. Here, because of (2) and (3), the percentage improvement in MSE of our model CARE, in comparison to the $M=3$ ensembles method [18], is $43.1\%$.
IV-B Further evaluation of CARE based on threshold $\zeta$, Table I
Furthermore, for the evaluation of CARE, we also examine in Table I the MSE $20\%$ or $10\%$ which is a threshold, i.e. $\zeta$, to detect instances where the model simply does not know the correct result (i.e. pixel-wise regression value) from the available input data, e.g. due to lack of spectral information or resolution. In such cases, models might choose to abstain from providing an answer and should be able to output “None of the above” for the result value of segmentation. A building density of $20\%$ with an estimated error of ±$\pm 10\%$ leads to the predicted density of $20\%$ being not useful, i.e. not accurate. In contrast, a building density of $80\%$ with an estimated error of ±$\pm 1\%$ leads to the predicted density of $80\%$ being useful and accurate.
@@ IV-B Further evaluation of CARE based on threshold $\zeta$, Table I
A prediction of the model with a high confidence indicates high reliability and trust for this particular prediction. Also, the features for high building density might not be clear in multi-spectral optical EO data. This is due to epistemic uncertainty, as it is induced by the lack of detail in the measurement. This is why a plausible output set that is an Open Set, is needed. Epistemic uncertainty is systematic, caused by lack of knowledge, and can be reduced by learning the characteristics of the quantity (e.g., high building density) using additional information (for example, in-situ measurements). Moreover, aleatoric uncertainty is statistical and related to randomness, and the sample not being a typical example of the quantity (e.g., very high or low building density).
@@
IV-C Evaluation of CARE using n-shot experiments in Table II
We evaluate the proposed model CARE in Table II using $n$-sample tests, for $n$ from $50$ to $10000$ samples. We examine both the MSE for the accuracy of the regression, i.e. (3) in (2), and the average error for the correctness of the confidence metric, i.e. $c_{i}$ in (4). Here, we observe that for $n=7500$ samples, the MSE is $0.00316$, while for $n=500$, the MSE is $0.0043$, i.e. the performance of the model improves when we have more labeled data. In addition, for the accuracy of the confidence metric of our model, the median error in Table II is $0.00190$ for $n=7500$ samples, while the median error is $0.0088$ for $n=500$. The performance of the model CARE, for confidence quantification, improves when we have more labeled data, i.e. $n=7500$ samples compared to $n=500$.
@@ IV-C Evaluation of CARE using n-shot experiments in Table II
For the evaluation of our model CARE, we also examine in Table II the MSE $20\%$ or $10\%$, similar to in Table I. We observe that the performance of the proposed model CARE for estimating the building density, i.e. how close the buildings are in the Sentinel-2 multi-spectral images, improves when the model has the ability to abstain in cases of low confidence. Moreover, Table II shows that when assuming that $10\%$ provide low confidence and our model should abstain, CARE gives improved performance, compared to the assumption of $20\%$. These results demonstrate the effectiveness of our algorithm in providing a reliable confidence metric for regression that leads to trustworthy and more accurate model predictions.
@@
IV-D Baseline model n-shot evaluation, Table III
In Table III, we evaluate the model Error-Sorting Confidence [6] on the Sentinel-2 PhilEO Bench building density estimation downstream task [1], to compare these results with the results of our proposed model CARE in Table II. Error-Sorting Confidence is used in Table III in the fine-tuning stage, i.e. when starting from and using a trained EO FM [1, 3]. Here, Error-Sorting is also an ablation study, i.e. using only the second term in (4). We observe that for $n=100$ training samples per region, the percentage improvement of CARE compared to Error-Sorting Confidence is $8.93\%$ for the MSE, and $73.47\%$ for the median error.
IV-E Qualitative evaluation of CARE
For the evaluation of CARE, we also examine qualitative results, in Figs. 2-4. In addition to the numerical evaluation metrics presented in the preceding subsections, we also evaluate the proposed model CARE qualitatively. For the error in the prediction, i.e. in (e), the variation in the regression output is, for example, because we underestimate in (b) the building density. In (f), the predicted uncertainty is $(1-\text{confidence})$, i.e. using (d). In the colorbar in (f), the maximum value is smaller than $1$, and this is desirable, i.e. we have less highlighted and emphasized yellow color.
@@ V Conclusion
In this paper, we have proposed the model CARE for pixel-wise regression tasks, for building density estimation from satellite Sentinel-2 data. Our model CARE computes and assigns a regression measure and a confidence metric to the output results of deep neural networks that solve regression problems. CARE not only predicts building density effectively but also provides reliable confidence scores that indicate trust in its predictions. The results reveal that CARE achieves a percentage improvement over Gaussian-Output NLL and other confidence estimation models. Evaluation metrics include the MSE, the average absolute error and the median absolute error, allowing comprehensive performance analysis against several baseline models. The experiments demonstrate that CARE significantly outperforms other baseline methods, showing improvements in various metrics. CARE shows a notable reduction in prediction error and demonstrates a strong correlation between confidence values and prediction errors. The results summarize the innovation introduced by the proposed model CARE, underscoring its dual capability of providing both accurate regression outputs and meaningful confidence metrics. We have shown that confidence quantification and assessment is critical for the deployment of models in real-life applications, particularly in EO tasks where uncertainties can have significant implications. By addressing the gap in confidence estimation for regression tasks in EO Foundation Models for building density estimation, this work contributes a valuable methodology and insights that can further enhance the practical application of neural networks in geospatial analysis. This underscores the importance of developing models that not only provide predictions but also quantify certainty, facilitating informed decision-making in critical areas such as urban planning and environmental monitoring.